The Linux HOWTO Index

by Greg Hankins, gregh@sunsite.unc.edu v2.10.41, 14 October 1997 This document contains an index to the Linux HOWTOs and mini-HOWTOs, as well as other information about the HOWTO project.

Contents

1 What Are Linux HOWTOs? 1

2 Where Do I Get Linux HOWTOs? 2 2.1 HOWTO Translations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

3 Index 3 3.1 HOWTOs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 3.2 mini-HOWTOs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 3.3 Special HOWTOs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 3.4 Unmaintained HOWTOs and mini-HOWTOs . . . . . . . . . . . . . . . . . . . . . . . . . . . 13

4 Writing and Submitting a HOWTO 13

5 Copyright 15

1 What Are Linux HOWTOs?

Linux HOWTOs are documents which describe in detail a certain aspect of configuring or using Linux. For example, there is the Installation HOWTO, which gives instructions on installing Linux, and the Mail HOWTO, which describes how to set up and configure mail under Linux. Other examples include the NET-3 HOWTO and the Printing HOWTO. HOWTOs are comprehensive docs - much like an FAQ but generally not in question-and-answer format. However, many HOWTOs contain an FAQ section at the end. There are several HOWTO formats available: plain text, PostScript, DVI, and HTML. In addition to the HOWTOs, there are a multitude of mini-HOWTOs on short, specific subjects. They are only available in plain text format.

2 Where Do I Get Linux HOWTOs?

HOWTOs can be retrieved via anonymous FTP from the following sites:

   • ftp://sunsite.unc.edu/pub/Linux/docs/HOWTO
   • ftp://tsx-11.mit.edu/pub/linux/docs/HOWTO

as well as many mirror sites. You can also browse HOWTOs in HTML format <http://sunsite.unc.edu/LDP/HOWTO/>. Many mirror

sites <http://sunsite.unc.edu/LDP/hmirrors.html> mirror the HTML files.

sunsite.unc.edu is heavily used, so please use a mirror site if possible. HOWTOs are also posted towards the beginning of the month to the Usenet newsgroup comp.os.linux.answers. There is a tool called NewstoHOWTO

<ftp://sunsite.unc.edu/pub/Linux/system/news/misc/> that will assemble the postings.

2.1 HOWTO Translations

HOWTO translations are available on sunsite.unc.edu

<ftp://sunsite.unc.edu/pub/Linux/docs/HOWTO/translations> and mirrors around the world. So far

there are:

• French (fr)

• German (de)

• Italian (it)

• Japanese (jp)

• Polish (pl)

• Spanish (es)

• Swedish (sv)

• Turkish (tr) If you know of any other translation projects, please let me know and I will add them to this list. If you are interested in getting your translations archived on sunsite.unc.edu, please read the directory structure specification at http://sunsite.unc.edu/pub/Linux/docs/HOWTO/translations/Directory-Structure and get in touch with me.

3 Index

3.1 HOWTOs

The following Linux HOWTOs are currently available:

• 3Dfx HOWTO, by Bernd Kreimeier <bk@gamers.org>. How to use 3Dfx graphics accelerator chip support. Updated 12 July 1997.

• AX25 HOWTO, by Terry Dawson <terry@perf.no.itg.telecom.com.au>. How to configure AX25 networking for Linux. Updated 20 June 1997.

• Access HOWTO, by Michael De La Rue <access-howto@ed.ac.uk>. How to use adaptive technology with Linux. Updated 28 March 1997.

• Alpha HOWTO, by David Mosberger <davidm@azstarnet.com>. Overview of Alpha systems and processors. Updated 6 June 1997.

• Assembly HOWTO, by François-René Rideau <rideau@ens.fr>. Information on programming in x86 assembly. Updated 7 September 1997.

• Benchmarking HOWTO, by André D. Balsa <andrewbalsa@usa.net>. How to do basic benchmarking. Updated 15 August 1997.

• BootPrompt HOWTO, by Paul Gortmaker <gpg109@rsphy1.anu.edu.au>. List of boot time arguments and overview of booting software. Updated 6 December 1996.

• Bootdisk HOWTO, by Graham Chapman <grahamc@zeta.org.au>. How to create a boot/root maintenance disk for Linux. Updated 4 April 1997.

• CD Writing HOWTO, by Winfried Trümper <winni@xpilot.org>. How to write CDs. Updated 10 August 1997.

• CDROM HOWTO, by Jeff Tranter <jeff tranter@pobox.com>. Information on CDROM drive compatibility for Linux. Updated 4 August 1997.

• Chinese HOWTO, by Hong Zhang <hongz@seas.ucla.edu>. How to configure Linux for use with the Chinese characterset. Updated 2 April 1997.

• Commercial HOWTO, by Martin Michlmayr <tbm@cyrius.com>. Listing of commercial software products for Linux. Updated 12 October 1997.

• Consultants HOWTO, by Martin Michlmayr <tbm@cyrius.com>. Listing of Linux consultants. Updated 12 October 1997.

• Cyrillic HOWTO, by Alexander L. Belikoff <abel@bfr.co.il>. How to configure Linux for use with the Cyrillic characterset. Updated 13 August 1997.

• DNS HOWTO, by Nicolai Langfeldt <janl@math.uio.no>. How to set up DNS. Updated 3 June • DOS to Linux HOWTO, by Guido Gonzato <guido@ibogfs.cineca.it>. How to move from DOS to

Linux. Updated 11 December 1996.

• DOSEMU HOWTO, by Uwe Bonnes <bon@elektron.ikp.physik.th-darmstadt.de>. HOWTO about the Linux MS-DOS Emulator, DOSEMU. Updated dosemu-0.64.4 (in progress), 15 March 1997.

• Danish HOWTO, by Niels Kristian Bech Jensen <nkbj@image.dk>. How to configure Linux for use with the Danish characterset. Updated 3 October 1997.

• Database HOWTO, by Al Dev (Alavoor Vasudevan) <deepav@pol.net>. How to setup PostgreSQL as a database server. Updated 19 July 1997.

• Disk HOWTO, by Stein Gjoen <sgjoen@nyx.net>. How to setup hard disk drives. Updated 12 August 1997.

• Distribution HOWTO, by Eric S. Raymond <esr@snark.thyrsus.com>. A list of Linux distributions. Updated 6 August 1997.

• ELF HOWTO, by Daniel Barlow <daniel.barlow@linux.org>. How to install and migrate to the ELF binary file format. Updated 14 July 1996.

• Emacspeak HOWTO, by Jim Van Zandt <jrv@vanzandt.mv.com>. How to use ’emacspeak’ with Linux. Updated 7 October 1997.

• Ethernet HOWTO, by Paul Gortmaker <gpg109@rsphy1.anu.edu.au>. Information on Ethernet hardware compatibility for Linux. Updated 2 February 1997.

• Finnish HOWTO, by Pekka Taipale <pjt@iki.fi>. How to configure Linux for use with the Finnish characterset. Updated 14 February 1996.

• Firewall HOWTO, by Mark Grennan, <markg@netplus.net>. How to set up a firewall using Linux. Updated 8 November 1996.

• French HOWTO, by Guylhem Aznar <guylhem@danmark.linux.eu.org>. How to configure Linux for use with the French characterset. Updated August 1997.

• Ftape HOWTO, by Kevin Johnson <kjj@pobox.com>. Information on ftape drive compatibility with Linux. Updated 15 March 1997.

• GCC HOWTO, by Daniel Barlow <daniel.barlow@linux.org>. How to set up the GNU C compiler and development libraries. Updated 28 February 1996.

• German HOWTO, by Winfried Trümper <winni@xpilot.org>. Information on using Linux with German-specific features. Updated 19 March 1997.

• Glibc2 HOWTO, by Eric Green <thrytis@imaxx.net>. How to install and migrate to the glibc2 library. Updated 31 August 1997.

• HAM HOWTO, by Terry Dawson <terry@perf.no.itg.telecom.com.au>. HOWTO configure amateur radio software for Linux. Updated 1 April 1997.

• HOWTO Index, by Greg Hankins <gregh@sunsite.unc.edu>. Index of HOWTO documents about Linux. Updated 12 October 1997.

• Hardware Compatibility HOWTO, by Patrick Reijnen <patrickr@bart.nl>. A list of hardware known to work with Linux. Updated 8 September 1997.

• Hebrew HOWTO, by Yair G. Rajwan <yair@hobbes.jct.ac.il>. How to configure Linux for use with the Hebrew characterset. Updated 12 September 1995.

• INFO-SHEET, by Michael K. Johnson <johnsonm@redhat.com>. Generic introduction to the Linux operating system. Updated 13 January 1997.

• IPX HOWTO, by Terry Dawson <terry@perf.no.itg.telecom.com.au>. How to install and configure IPX networking. Updated 29 March 1997.

• ISP Hookup HOWTO, by Egil Kvaleberg <egil@kvaleberg.no>. Basic introduction to hooking up to an ISP. Updated 14 February 1997.

• Installation HOWTO, by Eric S. Raymond <esr@snark.thyrsus.com>. How to obtain and install Linux. Updated 22 September 1997.

• Intranet Server HOWTO, by Pramod Karnad <karnad@indiamail.com>. How to setup a Linux Intranet server. Updated 7 August 1997.

• Italian HOWTO, by Marco “Gaio” Gaiarin <gaio@dei.unipd.it>. How to configure Linux for use with the Italian characterset. Updated 6 August 1997.

• Java-CGI HOWTO, by David H. Silber <dhs@orbits.com>. How to set up Java-capable CGI bin. Updated 18 November 1996.

• Kernel HOWTO, by Brian Ward <ward@blah.math.tu-graz.ac.at>. Upgrading and compiling the Linux kernel. Updated 26 May 1997.

• Keyboard and Console HOWTO, by Andries Brouwer <aeb@cwi.nl>. Information about the Linux keyboard, console and non-ASCII characters. Updated 28 August 1997.

• META-FAQ, by Michael K. Johnson <johnsonm@redhat.com>. A listing of Linux sources of information. Updated 20 December 1996.

• MGR HOWTO, by Vincent Broman <broman@nosc.mil>. Information on the MGR graphics interface for Linux. Updated 30 May 1996.

• MILO HOWTO, by David A. Rusling <david.rusling@reo.mts.dec.com>. How to use the Alpha Linux Miniloader (MILO). Updated 6th December 1996.

• Electronic Mail HOWTO, by Vince Skahan <vince@halcyon.com>. Information on Linux-based mail servers and clients. Updated 29 November 1995.

• NET-3 HOWTO, by Terry Dawson <terry@perf.no.itg.telecom.com.au>. Howto configure TCP/IP networking under Linux. Updated 20 August 1997.

• NFS HOWTO, by Nicolai Langfeldt <janl@math.uio.no>. How to set up NFS clients and servers. Updated 9 March 1997.

• NIS HOWTO, by Erwin Embsen <erwin@nioz.nl>. Information on using NIS/YP on Linux systems. Updated 14 November 1996.

• News HOWTO, by Vince Skahan <vince@halcyon.com>. Information on USENET news server and client software for Linux. Updated 29 November 1995.

• Optical Disk HOWTO, by Skip Rye <abr@brspc 0064.msd.ray.com>. How to use optical disk drives with Linux. Updated 17 July 1997.

• PCI HOWTO, by Michael Will <Michael.Will@student.uni-tuebingen.de>. Information on PCI-architecture compatibility with Linux. Updated 30 March 1997.

• PCMCIA HOWTO, by Dave Hinds <dhinds@allegro.stanford.edu>. How to install and use PCMCIA Card Services. Updated 9 October 1997.

• PPP HOWTO, by Robert Hart <hartr@interweft.com.au>. Information on using PPP networking with Linux. Updated 31 March 1997.

• Pilot HOWTO, by David H. Silber <pilot@orbits.com>. How to use a USR Pilot PDA with Linux. Updated 17 August 1997.

• Polish HOWTO, by Sergiusz Pawlowicz <ser@arch.pwr.wroc.pl>. Information on using Linux with Polish-specific features. Updated 5 January 1997.

• Printing HOWTO, by Grant Taylor <gtaylor+pht@picante.com>. HOWTO on printing software for Linux. Updated 23 September 1997.

• Printing Usage HOWTO, by Mark Komarinski <markk@auratek.com>. How to use the printing system for a variety of file types and options. Updated 24 March 1997.

• RPM HOWTO, by Donnie Barnes <djb@redhat.com>. How to use the Red Hat Package Manager (.rpm). Updated 8 April 1997.

• Reading List HOWTO, by Eric S. Raymond <esr@thyrsus.com>. Interesting books pertaining to Linux subjects. Updated 12 August 1997.

• SCSI HOWTO, by Drew Eckhardt <drew@PoohSticks.ORG>. Information on SCSI driver compatibility with Linux. Updated 30 August 1996.

• SCSI Programming HOWTO, by Heiko Eissfeldt <heiko@colossus.escape.de>. Information on programming the generic Linux SCSI interface. Updated 7 May 1996.

• SMB HOWTO, by David Wood <dwood@plugged.net.au>. How to use the Session Message Block (SMB) protocol with Linux. Updated 10 August 1996.

• SRM HOWTO, by David Mosberger <davidm@azstarnet.com>. How to boot Linux/Alpha using the SRM firmware. Updated 17 August 1996.

• Serial HOWTO, by Greg Hankins <greg.hankins@cc.gatech.edu>. How on use serial devices (modems, terminals) with Linux. Updated 6 July 1997.

• Serial Programming HOWTO, by Peter H. Baumann <Peter.Baumann@dlr.de>. How to use serial ports in programs. Updated 14 June 1997.

• Shadow Password HOWTO, by Michael H. Jackson <mhjack@tscnet.com>. How to obtain, install, and configure shadow passwords. Updated 3 April 1996.

• Slovenian HOWTO, by Primoz Peterlin <primoz.peterlin@biofiz.mf.uni-lj.si>. Information on using Linux with Slovenian-specific features. Updated 30 October 1996.

• Sound HOWTO, by Jeff Tranter <jeff tranter@pobox.com>. Sound hardware and software for the Linux operating system. Updated 4 August 1997.

• Sound Playing HOWTO, by Yoo C. Chung <wacko@laplace.snu.ac.kr>. How to play various sound formats under Linux. Updated 13 May 1997.

• Spanish HOWTO, by Gonzalo Garcia Agullo <Gonzalo.Garcia-Agullo@jrc.es>. Information on using Linux with Spanish-specific features. Updated 20 August 1996.

• teTeX HOWTO, by Robert Kiesling <kiesling@terracom.net>. How to install the teTeX package (TeX and LaTeX) under Linux. Updated 21 August 1997.

• Thai HOWTO, by Poonlap Veeratanabutr <poon-v@fedu.uec.ac.jp>. How to configure Linux for use with the Thai characterset. Updated 16 July 1997.

• Tips HOWTO, by Paul Anderson <paul@geeky1.ebtech.net>. HOWTO on miscellaneous tips and tricks for Linux. Updated 7 January 1997.

• UMSDOS HOWTO, by Jacques Gelinas <jacques@solucorp.qc.ca>. How to install and use the UMSDOS filesystem. Updated 13 November 1995.

• UPS HOWTO, by Harvey J. Stein <abel@netvision.net.il>. Information on using a UPS power supply with Linux. Updated 31 March 1997.

• UUCP HOWTO, by Vince Skahan <vince@halcyon.com>. Information on UUCP software for Linux. Updated 29 November 1995.

• User Group HOWTO, by Kendall Grant Clark <kclark@ntlug.org>. Tips on founding, maintaining, and growing a Linux User Group. Updated 9 October 1997.

• VAR HOWTO, by Martin Michlmayr <tbm@cyrius.com>. Listing of Linux value added resellers. Updated 12 October 1997.

• VMS to Linux HOWTO, by Guido Gonzato <guido@ibogfs.cineca.it>. How to move from VMS to Linux. Updated 18 March 1997.

• XFree86 HOWTO, by Eric S. Raymond <esr@snark.thyrsus.com>. How to obtain, install, and configure XFree86 3.2 (X11R6). Updated 5 September 1997.

• XFree86 Video Timings HOWTO, by Eric S. Raymond <esr@snark.thyrsus.com>. How to compose a mode line for XFree86. Updated 8 August 1997.

3.2 mini-HOWTOs

The following mini-HOWTOs are available (plain text only):

• 3 Button Mouse mini-HOWTO, by Geoff Short <geoff@kipper.york.ac.uk>. How to configure your mouse to use 3 buttons. Updated 11 July 1997.

• ADSM Backup mini-HOWTO, by Thomas Koenig <Thomas.Koenig@ciw.uni-karlsruhe.de>. How to install and use the ADSM backup program. Updated 15 January 1997.

• AI-Alife mini-HOWTO, by John A. Eikenberry <jae@ai.uga.edu>. Information about AI software for Linux. Updated 27 September 1997.

• Advocacy mini-HOWTO, by Paul L. Rogers <Paul.L.Rogers@li.org>. Suggestions on how to advocate the use of Linux. Updated 6 February 1997.

• Backup with MSDOS mini-HOWTO, by Christopher Neufeld <neufeld@physics.utoronto.ca>. How to backup Linux machines with MSDOS. Updated 5 August 1997.

• Battery Powered mini-HOWTO, by Hanno Mueller <hmueller@kabel.de>. How to reduce a Linux system’s power consumption. Updated 5 May 1997.

• Boca mini-HOWTO, by David H Dennis <david@freelink.net>. How to install a Boca 16-port serial card (Boca 2016). Updated 1 August 1997.

• BogoMips mini-HOWTO, by Wim C.A. van Dorst <baron@clifton.hobby.nl>. Information about BogoMips. Updated 26 August 1997.

• Bridge mini-HOWTO, by Chris Cole <chris@polymer.uakron.edu>. How to setup and ethernet bridge. Updated 23 August 1996.

• Bridge+Firewall mini-HOWTO, by Peter Breuer <ptb@it.uc3m.es>. How to setup and ethernet bridge and firewall. Updated 10 July 1997.

• Clock mini-HOWTO, by Ron Bean <rbean@execpc.com>. How to set and keep your clock on time. Updated December 1996.

• Colour ls mini-HOWTO, by Thorbjoern Ravn Andersen <ravn@dit.ou.dk>. How to set up the colours with ’ls’. Updated 7 August 1997.

• Comeau C++ mini-HOWTO, by Mark Swanson <ag010@freenet.carleton.ca>. How to set up Comeau C++. Updated 16 May 1997.

• DHCPcd mini-HOWTO, by Vladimir Vuksan <vuksan@veus.hr>. How to setup a DHCP client (dhcpcd). Updated 11 October 1997.

• DHCPd mini-HOWTO, by Paul Makeev <mac@RoSprint.net>. How to get Linux to serve DHCP information. Updated 24 June 1997.

• Dial On Demand mini-HOWTO, by Rodney van den Oever <Rodney.van.den.Oever@tip.nl>. How to use dial-on-demand dialing to dial an ISP. Updated 29 February 1997.

• Diald mini-HOWTO, by Harish Pillay <h.pillay@ieee.org>. How to use ’diald’ to dial an ISP. Updated 3 June 1996.

• Diskless mini-HOWTO, by Robert Nemkin <buci@math.klte.hu>. How to set up a diskless Linux box. Updated 31 May 1996.

• Dynamic IP Hacks mini-HOWTO, by Michael Driscoll <fenris@lightspeed.net>. Tips and tricks for people with dynamic IP addresses. Updated 25 May 1997.

• Ext2fs Undeletion mini-HOWTO, by Aaron Crane <aaronc@pobox.com>. How to retrieve deleted files from an ext2 filesystem. Updated 4 August 1997.

• Fax Server mini-HOWTO, by Erez Strauss <erez@newplaces.com>. How to setup a fax server. Update 24 September 1997.

• GIS-GRASS mini-HOWTO, by David A. Hastings <dah@ngdc.noaa.gov>. How to install Geographic Information System (GIS) software. Updated 15 July 1997.

• GTEK BBS-550 mini-HOWTO, by Wajihuddin Ahmed <wahmed@sdnpk.undp.org>. How to setup the GTEK BBS-550 multiport board with Linux. Updated 20 August 1997.

• Gravis Ultra Sound mini-HOWTO, by Nicolai Langfeldt <janl@math.uio.no>. How to install the Gravis Ultra Sound PnP card. Updated 14 September 1997.

• IO Port Programming mini-HOWTO, by Riku Saikkonen <Riku.Saikkonen@hut.fi>. How to use I/O ports in C programs. Updated 30 March 1997.

• IP Alias mini-HOWTO, by Harish Pillay <h.pillay@ieee.org>. How to use IP aliasing. Updated 13 January 1997.

• IP Masquerade mini-HOWTO, by Ambrose Au <achau@wwonline.com>. How to use IP masquerading. Updated 1 January 1997.

• IP Subnetworking mini-HOWTO, by Robert Hart <hartr@interweft.com.au>. Why and how to subnetwork an IP network. Updated 31 March 1997.

• IP Connectivity mini-HOWTO, by Michael Strates <Michael Strates@apana.org.au>. How to get mail and news over a dialup connection. Updated 30 August 1997.

• Jaz Drive mini-HOWTO, by Bob Willmot <bwillmot@cnct.com>. How to use an Iomega Jaz drive with Linux. Updated 4 June 1997.

• Kerneld mini-HOWTO, by Henrik Storner <storner@osiris.ping.dk>. How to use ’kerneld’ (dynamic module loading). Updated 19 July 1997.

• LBX mini-HOWTO, by Paul D. Smith <psmith@BayNetworks.com>. How to use Low-Bandwidth X (LBX). Updated 17 July 1997.

• Large Disk mini-HOWTO, Andries Brouwer <aeb@cwi.nl>. How to use disks with > 1024 cylinders. Updated 26 July 1996.

• Linux+DOS+Win95 mini-HOWTO, by Alan L. Wendt <alan@ez0.ezlink.com>. How to use Linux and DOS and Windows95 together. Updated 10 September 1996.

• Linux+DOS+Win95+OS2 mini-HOWTO, by Mike Harlan <r3mdh@imperium.net>. How to use Linux and DOS and OS/2 and Win95 together. Updated 23 July 1997.

• Linux+OS2+DOS mini-HOWTO, by Hamish Moffatt <moffatt@yallara.cs.rmit.edu.au>. How to use Linux and OS/2 and DOS together. Updated 20 May 1996.

• Linux+FreeBSD mini-HOWTO, by Niels Kristian Bech Jensen <nkbj@image.dk> How to use Linux and FreeBSD together. Updated 3 October 1997.

• Linux+NT-Loader mini-HOWTO, by Bernd Reichert <reichert@dial.eunet.ch>. How to use Linux and the Windows NT boot loader together. Updated 2 September 1997.

• Linux+Win95 mini-HOWTO, by Jonathan Katz <jkatz@in.net>. How to use Linux and Windows95 together. Updated 25 June 1996.

• Loadlin+Win95 mini-HOWTO, by Chris Fischer <protek@brigadoon.com>. How to use Linux and Windows95 together, using loadlin. Updated 15 September 1997.

• Locales mini-HOWTO, by Peeter Joot <peeter joot@vnet.ibm.com>. How to configure Linux to use locales. Updated 21 July 1997.

• MIDI+SB mini-HOWTO, by Hideki Saito <hideki@eskimo.com>. How to a MIDI keyboard with a Sound Blaster. Updated 13 October 1996.

• Mail Queue mini-HOWTO, by Leif Erlingsson <Leif.Erlingsson@leif@lege.com>. How to queue remote mail and deliver local mail. Updated 3 September 1997.

• Mail2News mini-HOWTO, by Robert Hart <iweft@ipax.com.au>. How to set up a mail to news gateway. Updated 4 November 1996.

• Man Page mini-HOWTO, by Jens Schweikhardt <schweikh@noc.dfn.de>. How to write man pages. Updated August 1997.

• Modules mini-HOWTO, by Riley H. Williams <rhw@bigfoot.com>. How to setup and configure kernel modules. Updated 15 August 1997.

• NFS-Root mini-HOWTO, by Andreas Kostyrka <andreas@ag.or.at>. How to set up diskless Linux machines. Updated 8 August 1997.

• NFS-Root-Client mini-HOWTO, by Ofer Maor <ofer@hadar.co.il>. How to set up diskless Linux machines using NFS. Updated 1 July 1997.

• Netscape+Proxy mini-HOWTO, by Sarma Seetamraju <sarma@usa.net>. How to set up a proxy server for Netscape. Updated 15 August 1997.

• Offline Mailing mini-HOWTO, by Gunther Voet <gunther.voet@ibm.net>. How to set up email addresses without a dedicated Internet connection. Updated 15 September 1997.

• Online Support mini-HOWTO, by lilo <TaRDiS@mail.utexas.edu>. Information about the Linux Internet Support Cooperative. Updated 7 May 1996.

• PLIP mini-HOWTO, by Andrea Controzzi <controzz@cli.di.unipi.it>. How to set up PLIP (Parallel Line Interface Protocol). Updated 22 August 1996.

• PPP-over-minicom mini-HOWTO, by Winfried Trümper <winni@xpilot.org>. How to use minicom to start PPP. Updated 9 December 1996.

• Pager mini-HOWTO, by Chris Snell <chris@unm.dorm.net>. How to set up an alphanumeric paging gateway. Updated 29 August 1997.

• Partition mini-HOWTO, by Kristian Koehntopp <kris@koehntopp.de>. How to choose disk partitions. Updated 15 July 1997.

• Print2Win mini-HOWTO, by Harish Pillay <h.pillay@ieee.org>. How to use SAMBA to print to Windows. Updated 3 June 1996.

• Process Accounting mini-HOWTO, by Albert M.C. Tam <bertie@scn.org>. How to set up process accounting. Updated 8 August 1997.

• Proxy ARP mini-HOWTO, by Al Longyear <longyear@netcom.com>. A small treatise on the use of Proxy ARP. Updated 5 December 1994.

• Proxy ARP Subnet mini-HOWTO, by Bob Edwards <Robert.Edwards@anu.edu.au>. How to use proxy ARP with subnetting. Updated August 1997.

• Public Web Browser mini-HOWTO, by Donald B. Marti Jr. <dmarti@agdia.com>. How to set up a guest account to use a WWW browser. Updated 2 February 1997.

• Qmail+MH mini-HOWTO, by Christopher Richardson <rdn@tara.n.eunet.de>. How to install qmail and MH. Updated 13 June 1997.

• Quota mini-HOWTO, by Albert M.C. Tam <bertie@scn.org>. How to set up disk quotas. Updated 8 August 1997.

• RCS mini-HOWTO, by Robert Kiesling <kiesling@terracom.net>. How to use RCS (Revision Control System). Updated 14 August 1997.

• RPM+Slackware mini-HOWTO, by Dave Whitinger, <wolf@redhat.com>. How to install the Red Hat Package Manager (RPM) under Slackware. Updated 29 August 1997.

• Remote Boot mini-HOWTO, by Marc Vuilleumier Stückelberg <Marc.VuilleumierStuckelberg@cui.unige.ch>. How to set up a server-based boot selector. Updated August 1997.

• Remote X Apps mini-HOWTO, by Vincent Zweije <zweije@xs4all.nl>. How to run remote X applications. Updated 5 August 1997.

• SLIP+proxyARP mini-HOWTO, by Dave Kennedy <davek@muscle.net>. How to use SLIP and Proxy ARP to connect LANs to the net. Updated 20 November 1996.

• SLIP-PPP Emulator mini-HOWTO, by Irish <irish@eskimo.com>. How to use SLIP-PPP emulators with Linux. Updated 7 August 1997.

• Sendmail+UUCP mini-HOWTO, by Jamal Hadi Salim <jamal@glcom.com>. How to use sendmail and UUCP together. Updated 15 March 1997.

• Software Building mini-HOWTO, by Mendel Leo Cooper <thegrendel@theriver.com>. How to build software packages. Updated 13 August 1997.

• Software RAID mini-HOWTO, by Linas Vepstas <linas@fc.net>. How to configure software RAID. Updated 29 September 1997.

• Soundblaster-16 mini-HOWTO, by Eric Z. Ayers <Eric.Ayers@compgen.com>. How to install the Soundblaster 16 PnP card. Updated 14 January 1997.

• Soundblaster-AWE64 mini-HOWTO, by Bucky LaDieu <nega@vt.edu>. How to install the Soundblaster AWE64 PnP card. Updated 25 June 1997.

• StarOffice mini-HOWTO, by Matthew Borowski <mkb@poboxes.com>. Information on installing the StarOffice suite. Updated 7 July 1997.

• Swap Space mini-HOWTO, by H. Peter Anvin <hpa@yggdrasil.com>. How to share swap space between Linux and Windows. Updated 19 June 1995.

• Term Firewall mini-HOWTO, by Barak Pearlmutter <bap@cs.unm.edu>. How to use ’term’ over a firewall. Updated 15 July 1997.

• Token Ring mini-HOWTO, by Mike Eckhoff <meckhoff@wayne.esu1.k12.ne.us>. How to use token ring cards. Updated 30 April 1997.

• Upgrade mini-HOWTO, by Greg Louis <glouis@dynamicro.on.ca>. How to upgrade your Linux distribution. Updated 6 June 1996.

• VPN mini-HOWTO, by Árpád Magosányi <mag@bunuel.tii.matav.hu>. How to set up a VPN (Virtual Private Network). Updated 3 May 1997.

• Virtual wu-ftpd mini-HOWTO, by Winfried Trümper <winni@xpilot.org>. How to set up wu-ftpd with virtual domains. Updated 27 January 1997.

• Visual Bell mini-HOWTO, by Alessandro Rubini <alessandro.rubini@linux.it>. How to disable audible bells, and enable visual bells. Updated August 1996.

• Win95+Win+Linux mini-HOWTO, by Robert Goodwin <Robert.Goodwin@mcc.ac.uk>. How to get Windows95, Windows 3.x and Linux to work. Updated 4 August 1996.

• Windows Modem Sharing mini-HOWTO, by Friedemann Baitinger <baiti@herrenberg.netsurf.de>. How to setup Windows to use a shared modem on a Linux machine. Updated 12 June 1997.

• WordPerfect mini-HOWTO, by Wade Hampton <whampton@staffnet.com¿>. How to set up WordPerfect for Linux. Updated 13 August 1997.

• X Big Cursor mini-HOWTO, by Joerg Schneider <schneid@ira.uka.de>. How to use enlarged cursors with XWindows. Updated 11 August 1997.

• XFree86-XInside mini-HOWTO, by Marco Melgazzi <marco@techie.com>. How to convert XFree86 to XInside modelines. Updated September 1997.

• xterm Title mini-HOWTO, by Winfried Trümper <winni@xpilot.org>. How to put strings into the titlebar of an xterm. Updated 22 October 1996.

• ZIP Drive mini-HOWTO, by Grant Guenther <grant@torque.net>. How to use an Iomega ZIP drive with Linux. Updated 15 April 1996.

• ZIP Install mini-HOWTO, by John Wiggins <jwiggins@comp.uark.edu>. How to install Linux onto a ZIP drive. Updated 5 December 1996.

3.3 Special HOWTOs

The High Availability HOWTO, by Harald Milz <hm@seneca.muc.de> is available at

http://sunsite.unc.edu/pub/Linux/ALPHA/linux-ha/High-Availability-HOWTO.html. It is not

included with the HOWTO collection because it relies on figures and cannot be distributed in all supported formats.

The Graphics HOWTO, by Michael J. Hammel <mjhammel@csn.net> is available at

http://www.csn.net/ mjhammel/linux/lgh.html. It is not included with the HOWTO collection

because it needs to use a lot of images, which don’t translate to other output formats.

3.4 Unmaintained HOWTOs and mini-HOWTOs

There are a number of unmaintained documents at

ftp://sunsite.unc.edu/pub/Linux/docs/HOWTO/unmaintained. These are kept around since old

documentation is sometimes better than none. However, you should be aware that you are reading old documentation.

4 Writing and Submitting a HOWTO

If you are interested in writing a HOWTO or mini-HOWTO, please get in touch with me FIRST!!! at

gregh@sunsite.unc.edu.

Here are a few guidelines that you should follow when writing a HOWTO or mini-HOWTO:

• Try to use meaningful structure and organization, and write clearly. Remember that many of the people reading HOWTOs do not speak English as their first language.

• If you are writing a HOWTO, you must use the SGML-Tools package, available from http://www.pobox.com/ cg/sgmltools, to format the HOWTO. This package allows us to produce LaTeX (for DVI and PostScript), plain text, and HTML from a single source document, and was designed specifically for the HOWTOs. This also gives all of the HOWTOs a uniform look.

• If you are writing a mini-HOWTO, you can either use SGML (as described above) or HTML. Plain text submissions are no longer accepted! In the near future, I will start distributing mini-HOWTOs in plain text and HTML.

• Make sure that all of the information is correct. I can’t stress this enough. When in doubt, speculate, but make it clear that you’re only guessing.

• Make sure that you are covering the most recent version of the available software. Also, be sure to include full instructions on where software can be downloaded from (FTP site name, full pathname), and the current version number and release date of the software.

• Include an FAQ section at the end, if appropriate. Many HOWTO documents need an “FAQ” or “Common Problems” section to cover information which can’t be covered in the regular text.

• Use other HOWTOs or mini-HOWTOs as a model! The SGML source to the HOWTOs is available on Linux FTP sites.

• Make sure that your name, email address, date, and a version number is near the beginning of the document. You could also include WWW addresses and a snail mail address if you want. The standard header is:

        Title
        Author's name and email address
        Version number and date

for example:

        The Linux HOWTO Index
        by Greg Hankins
        v2.10.29, 31 July 1997

• Lastly, be prepared to receive questions and comments about your writing. There are several hundreds of accesses to the HOWTO collection every day from around the world!

After you have written the HOWTO, mail it to me. If you used SGML-Tools, simply mail me the SGML source; I take care of formatting the documents. I’ll also take care of archiving the HOWTOs on sunsite.unc.edu and posting them to the various newsgroups.

It is important that you go through me when submitting a HOWTO, as I maintain the archives and need to keep track of what HOWTOs are being written and who is doing what.

Then, all you have to do is send me periodic updates whenever appropriate.

5 Copyright

Unless otherwise stated, Linux HOWTO documents are copyrighted by their respective authors. Linux HOWTO documents may be reproduced and distributed in whole or in part, in any medium physical or electronic, as long as this copyright notice is retained on all copies. Commercial redistribution is allowed and encouraged; however, the author would like to be notified of any such distributions.

All translations, derivative works, or aggregate works incorporating any Linux HOWTO documents must be covered under this copyright notice. That is, you may not produce a derivative work from a HOWTO and impose additional restrictions on its distribution. Exceptions to these rules may be granted under certain conditions; please contact the Linux HOWTO coordinator at the address given below.

In short, we wish to promote dissemination of this information through as many channels as possible. However, we do wish to retain copyright on the HOWTO documents, and would like to be notified of any plans to redistribute the HOWTOs.

If you have questions, please contact Greg Hankins, the Linux HOWTO coordinator, at

gregh@sunsite.unc.edu via email.